# Prediction Sets and Conformal Inference with

### v2

@@ 5 Numerical experiments
  • Model A. The unobserved outcomes $Y_{i}$ are generated according to a mixture model with heteroscedasticity, similar to the simulations in lei2014DistributionfreePrediction.
@@
  • Model B. The unobserved outcomes are generated in the same way as in Equation 12 above. We randomly select 20% of the unobserved outcomes and they are interval censored and in this case and $(Y^{L}_{i},Y_{i}^{U})$ have a discrete distribution , while the remaining 80% are observed without censoring. Let $\iota_{i}$ be a random draw from a Bernoulli distribution with probability 0.2,
@@
    This model is designed to evaluate the performance of the conformalized prediction set under fixed censoring. We choose to censor 20% of the outcomes, as approximately 20% of the income data is censored in the dataset used for our empirical studies.
@@
    and $(Y_{i}^{L},Y_{i}^{U})$ are generated accordng to Equation 14. In this case, the conditional distribution is unimodal and skewed.
@@
The observations $\left\{(X_{i},Y_{i}^{L},Y_{i}^{U}),i=1,\dots,n\right\}$, are randomly split into a training set $\mathcal{I}_{1}$ and a calibration set $\mathcal{I}_{2}$, with the training set containing 75% of the observations. We estimate the prediction set $\hat{C}_{S}(x)$ based on the kernel smoothing estimator $P_{n}(C;x)$ with the Epanechnikov kernel for the conditional distribution $P(C;x)$, as described in Equation (8), using the observations in the training set. The conformal prediction sets $\tilde{C}_{S}$ are then constructed according to Equation (11) using the calibration set. Local conformal prediction sets $\tilde{C}_{S,\text{loc}}$ are obtained using the local conformalization procedure, where the support $\mathcal{X}$ is divided into 5 equal-sized bins.
@@
We then compare the average coverage and volume of the conformal prediction set $\tilde{C}$ and the local conformal prediction set $\tilde{C}_{\text{loc}}$ over 100 repetitions with two alternative conformal prediction intervals: $\tilde{C}_{q2}$, based on a quantile regression assuming a quadratic model, and $\tilde{C}_{q3}$, based on a quantile regression assuming a cubic model. See romano2019ConformalizedQuantile and Appendix E for more details about construction of prediction interval based on quantile methods. The coverage is computed by randomly generating 5,000 samples of $(Y^{L},Y^{U})$ and calculating the proportion of intervals covered by the prediction sets. The volume is calculated as the integrated Lebesgue measure of the prediction sets over the support $\mathcal{X}$.
@@
Figure 2 - 4 show the comparison of the prediction sets from the numerical experiments. The top two rows plot the observed sample $(Y_{i}^{L},Y_{i}^{U})$ against $X_{i}$ as well as the conformal prediction sets $C(X_{i})$. The last row shows the integrated volume of the prediction sets and the coverage property. In Figure 2 and Figure 3, the prediction sets $\tilde{C}_{S}$ and $\tilde{C}_{S,\text{loc}}$ capture the bifurcation region of the conditional distribution and have smaller volume compared to the quantile regression-based prediction sets. Even when the conditional distribution is uni-modal, and the oracle prediction set is a single interval for each $x\in\mathcal{X}$, our proposed estimators still have smaller volume as demonstrated in Figure 4. In terms of coverage, $\tilde{C}_{S},\tilde{C}_{S,\text{loc}}$ and $\tilde{C}_{q3}$ appear to have coverage close to the nominal leval $0.91-\alpha=$ 0.9. Notably, $\tilde{C}_{S,\text{loc}}$ appears to have better coverage than $\tilde{C}_{S}$ near the boundaries of the support $\mathcal{X}$.
(a) Conformal prediction set $\tilde{C}_{S}$
(b) Local conformal prediction set $\tilde{C}_{S,\text{loc}}$
(c) Conf. pred. quadratic QR
(d) Conf. pred. cubic QR
(e) Integrated volume
(f) Coverage
Figure 2: Comparison of prediction sets in simulation study. The first four sub-figures show the interval censored ourcomes and the conformal prediction sets. Subfigure (e) shows the integrated volume of the prediction sets, and subfigure (f) shows the coverage of the prediction sets.
(a) Conformal prediction set
(b) Local conformal prediction set
(c) Conf. pred. quadratic QR
(d) Conf. pred. cubic QR
(e) Integrated volume
(f) Coverage
Figure 3: Comparison of prediction sets in simulation study. The first four sub-figures show the interval censored ourcomes and the conformal prediction sets. Subfigure (e) shows the integrated volume of the prediction sets, and subfigure (f) shows the coverage of the prediction sets.
(a) Conformal prediction set
(b) Local conformal prediction set
(c) Conf. pred. quadratic QR
(d) Conf. pred. cubic QR
(e) Integrated volume
(f) Coverage
Figure 4: Comparison of prediction sets in simulation study. The first four sub-figures show the interval censored ourcomes and the conformal prediction sets. Subfigure (e) shows the integrated volume of the prediction sets, and subfigure (f) shows the coverage of the prediction sets.
@@ 6.1 US CPS data
We applied our method to the 2023 US Current Population Survey (CPS) dataset and construct prediction interval for their income given years of education and experience. Following heckman2008EarningsFunctions, we construct the following variables from the dataset for individuals who report being American citizen and employed:
@@
  • Education: Similar to heckman2008EarningsFunctions, we construct the years of education for each individual by converting the highest qualification they have obtained into years of education.
@@
The CPS dataset consists of 146,133 observations. There are 53,171 individuals who satisfy the demographic restrictions above. Among the individuals that satisfy the demographic criteria, 39,333 reported their income either exactly or as a band and the response rate is thus around 74%. The issue of nonresponse in surveys has been studied in the literature; see for example, lillard1986WhatWe and bollinger2019TroubleTails, where they document an increasing proportion of nonresponse of earnings in survey data and the potential bias it introduces. We will not address the nonresponse issue in this paper, but it is an important consideration in practice. Our proposed predictions could assist in the imputation of missing values in a robust manner; see lei2021ConformalInference. In fact, allowing respondents to report their income in bands could help mitigate the non-response problem. For the 39,333 individuals who reported their incomes and satisfy the demographic restrictions, 8,819 of them reported either their gross or net income as a band, which thus amounts to roughly 22.4%. of the respondents. Figure 5 shows the proportion of individuals who reported their income in a range by years of education.
@@
In our empirical study, we first randomly reserve 20% of the data as a hold-out dataset. The remaining 80% is then further divided into training and calibration sets, with 75% of the data used for training to estimate the prediction interval $\hat{C}_{I}$ and 25% used for constructing the conformalized prediction interval $\tilde{C}_{I}$. Table LABEL:tab:CPS presents the average estimated prediction intervals across the 100 iterations. The method column indicates the way we estimate the prediction interval. “P” stands for prediction interval $\hat{C}_{I}$ in (7) and “CP” stands for conformal prediction interval $\tilde{C}_{I}$. If “M” is appended, it indicates the prediction intervals are estimated with imputed income for individuals who reported their income in an interval. It can be seen from the table that we would in general predict a higher income for individuals with more years of education.
@@
The impact of including interval-censored income data on the prediction interval is more pronounced at smaller $\alpha$ values, as evidenced by the results for $\alpha=0.1$ in the table. This is because, with a 90% coverage target, the tails of the income distribution become more critical. Given that 20% of the observations are reported in ranges and many fall into the highest category, incorporating interval-censored data shifts the right tail of the income distribution. Consequently, the prediction intervals constructed with interval-censored data are significantly different from the ones constructed with imputed income data.
We also compare the coverage of the income intervals in the hold-out calibration set by prediction intervals constructed with interval-censored data and with imputed data. In Figure 6, we show the comparison of the coverages. The conformal prediction interval constructed with interval-censored data has a higher coverage at the chosen level of $0.91-\alpha=$ 0.9, while the prediction interval constructed with imputed data has a lower coverage. Table 2 below shows the coverage of the interval censored incomes in the hold-out dataset for the conformal prediction interval $\tilde{C}$ that are constructed using both interval-censored and exactly reported income data (shown in the “Censored” columns), as well as for the conformal prediction intervals constructed when the interval-censored incomes are imputed (shown in the “Imputed” columns). The significant difference in prediction intervals constructed with interval-censored data and with imputed data may be due to the fact that there are only five categories for individuals who report their income in a range. This seems somewhat coarse in terms of informing about the true income distribution. Notice that the conformal coverage using the censored model has coverage close to the nominal rate while the Imputed Coverage is almost always lower. Note that in Tables 1 an 2 when we display coverages of .5 and .1, this provides an empirical estimate of the 50% and 10% prediction sets, i.e., sets where respectively 50% and 90% of where the data are and should be informative about the shape of the data distribution. Notice that standard reporting of confidence intervals on parameters is meant to summarize sampling uncertainty, while here not only do these conformalized sets take into account sampling uncertainty, but also these are estimates of prediction sets of the underlying distributions of interest.
@@ 0.9 CP 10 (37,280 - 47,224) (51,494 - 64,217) (50,508 - 59,915) (64,290 - 77,790)
(a) Years of experience = 10
(b) Years of experience = 20
(c) Years of experience = 30
(d) Years of experience = 40
Figure 6: Comparison of coverage of prediction intervals constructed with interval-censored data and with imputed data for $\alpha=0.1$.
7 Conclusion
@@ A.1 Proofs for Section 2
which shrinks nicely to $(a,b)$ folland1999RealAnalysis, and as a reult for some $k^{*}$, $\delta=\int_{A_{k^{*}}}f_{Y^{L},Y^{U}}>0$. Let $P_{\pi/2}$ be the one defined in Lemma A.1 and $P_{\pi/2}^{\prime}$ be the positive measure constructed with the same method for the density function $f_{Y^{L},Y^{U}}\mathbb{I}\left\{A_{k^{*}}^{c}\right\}$. For any $y\in[a,b]$, and any $\lambda>0$, let $0<\varepsilon<\delta/\lambda$, and define the following probability distribution, for any interval $B\subset\mathbb{R}$,
@@
Since $P(Y\in A)\geq P([Y^{L},Y^{U}]\subset A)$, it is clear that the outer region of the level set with $1-\alpha$ covergage of $(Y^{L},Y^{U})$ also has $1-\alpha$ coverage for $Y$. But the outer region is in general not the smallest valid prediction set.
The following lemma shows that even though our approach is conservative in terms of coverage, its conservativeness is only due to partial identification, which is necessary without assumeing the mechanism of censoring.
@@
$P_{2}(A)=\int_{A_{\pi}}g_{x}\mathbb{I}\left\{A^{\prime}_{m}\right\},$
@@ A.2 Proofs for Section 3
In this subsection, we first prove the uniform consistency of the proposed prediction set when the oracle prediction set is an interval under random censoring. Then the result is extended to the case allowing for fixed censoring and when oracle prediciton set is a union of multiple intervals.
@@
    In the first case, we have $\hat{\tau}_{1}-\hat{\tau}_{0}\leq\tau_{1}(x)-\tau_{0}(x)$, if $E_{0,x}$ also holds, then $P(\hat{\tau}_{0},\hat{\tau}_{1};x)\leq 1-\alpha-\delta$ by Assumption 3.2, while $P_{n}(\hat{\tau}_{0},\hat{\tau}_{1};x)\geq 1-\alpha$, which implies that $\left\lvert P_{n}(\hat{\tau}_{0},\hat{\tau}_{1};x)-P\left(\hat{\tau}_{0},\hat{\tau}_{1};x\right)\right\rvert>\delta>\kappa$. Hence $E_{1,x}E_{0,x}\subset E^{c}$.
@@
For any $\varepsilon>0$, let $\delta$ be chosen as in Assumption 3.5. Let $E$ denote the event when $\sup_{C,x}\left\lvert P_{n}(C;x)-P(C;x)\right\rvert\leq\psi_{n}$. For sufficiently large $n$, $\psi_{n}\leq\frac{\delta}{2}$. For any $x\in\mathcal{X}$, if $\mu(\hat{C}(x)\bigtriangleup C^{*}(x))>\varepsilon$ and $\mu(\hat{C}(x))\leq\mu(C^{*}(x))$, then $P(\hat{C}(x);x)\leq 1-\alpha-\delta$, and $P_{n}(\hat{C}(x);x)\geq 1-\alpha-\psi_{n}$. Hence $\left\lvert P_{n}(\hat{C};x)-P(\hat{C};x)\right\rvert>\frac{\delta}{2}\geq\psi_{n}$. On the other hand, if $\mu(\hat{C}(x)\bigtriangleup C^{*}(x))>\varepsilon$ and $\mu(\hat{C}(x))>\mu(C^{*}(x))$, then $P_{n}(C(x);x)<1-\alpha-\psi_{n}$ while $P(C(x);x)\geq 1-\alpha$ and $\left\lvert P_{n}(C(x);x)-P(C(x);x)\right\rvert>\psi_{n}$. That is $P(\sup_{x}\mu(\hat{C}(x)\bigtriangleup C^{*}(x))>\varepsilon)\leq P(E^{c})\to 0$. ∎
@@ A.3 Proofs for results in Section 4
Let $s_{i}=s(Y^{L}_{i},Y^{U}_{i},X_{i};\hat{C})$, then $[Y^{L}_{n+1},Y^{U}_{n+1}]\subset\tilde{C}(X_{n+1})$ if and only if $s_{n+1}\leq\vartheta_{1-\alpha}$. Since $s_{i}$’s are exchangeable for $i\in\mathcal{I}_{2}$, $P\left(s_{n+1}\leq\vartheta_{1-\alpha}\right)\geq 1-\alpha$ by Lemma 2 in romano2019ConformalizedQuantile. ∎
@@
Fix an $\varepsilon>0$, on the event $\sup_{x}\mu\left(\hat{C}_{I}(x)\bigtriangleup C_{I}(x)\right)<\frac{\varepsilon}{2}$, we know that
$\left\lvert s(y^{l},y^{u},x;\hat{\tau}_{0},\hat{\tau}_{1})-s(y^{l},y^{u},x;\tau_{0},\tau_{1})\right\rvert<\frac{\varepsilon}{2},$
for all $y^{l},y^{u}$ and $x$. Let $s_{[j]},s^{*}_{[j]}$ be the corresponding order statistics for $s_{j}$ and $s^{*}_{j},j\in\mathcal{I}_{2}$. Let $s_{[l]}=s_{[l]}$ when $l$ is an integer and $s_{[l]}=s_{[\lfloor l\rfloor+1]}$ when $l$ is not an integer. We have $s_{[(n_{2}+1)(1-\alpha)]}>s^{*}_{[(n_{2}+1)(1-\alpha)]}-\varepsilon$ provided that $s^{*}_{[(n_{2}+1)(1-\alpha)]}-s^{*}_{[(n_{2}+1)(1-\alpha)-1]}<\frac{\varepsilon}{2}$ which happens with high probability for sufficiently large $n_{2}$. Similarly we have $s_{[(n_{2}+1)(1-\alpha)]}<s^{*}_{[(n_{2}+1)(1-\alpha)]}+\varepsilon$. Thanks to the consistency of sample quantile, we have that $s^{*}_{[(n_{2}+1)(1-\alpha)]}\to_{p}Q_{1-\alpha}(s^{*})$ as $n_{2}\to\infty$. By definition, $[y^{L},y^{U}]\subset[t_{0}(x),t_{1}(x)]$ if and only if $s(y^{L},y^{U},x;t_{0},t_{1})\leq 0$. Since $Q_{1-\alpha}(s^{*})=\inf\left\{z:P(s^{*}\leq z)\geq 1-\alpha\right\}$, combined with
$P(s^{*}\leq 0)=P\left(\tau_{0}\leq Y^{L}\leq Y^{U}\leq\tau_{1}\right)\geq 1-\alpha,$
and for any $z<0$, $P(s^{*}\leq z)<1-\alpha$ by Assumption 3.2, we have $Q_{1-\alpha}(s^{*})=0$. Hence $\vartheta_{1-\alpha}=o_{p}(1)$. ∎

### v3

@@ 5 Numerical experiments
  • Model A. The unobserved outcomes $Y_{i}$ are generated according to a mixture model with heteroscedasticity, similar to the simulations in Lei and Wasserman, (2014).
@@
  • Model B. The unobserved outcomes are generated in the same way as in Equation 14 above. We randomly select 20% of the unobserved outcomes and they are interval valued and in this case and $(Y^{L}_{i},Y_{i}^{U})$ have a discrete distribution , while the remaining 80% are observed without censoring. Let $\iota_{i}$ be a random draw from a Bernoulli distribution with probability 0.2,
@@
    This model is designed to evaluate the performance of the conformalized prediction set under fixed censoring. We choose to censor 20% of the outcomes, as approximately 20% of the income data is interval valued in the dataset used for our empirical studies.
@@
    and $(Y_{i}^{L},Y_{i}^{U})$ are generated according to Equation 16. In this case, the conditional distribution is unimodal and skewed.
@@
The observations $\left\{(X_{i},Y_{i}^{L},Y_{i}^{U}),i=1,\dots,n\right\}$, are randomly split into a training set $\mathcal{I}_{1}$ and a calibration set $\mathcal{I}_{2}$, with the training set containing 75% of the observations. We estimate the prediction set $\hat{C}(x)$ based on the kernel smoothing estimator $P_{n}(C;x)$ with the Epanechnikov kernel for the conditional distribution $P(C;x)$, as described in Equation (8), using the observations in the training set. The conformal prediction sets $\tilde{C}$ are then constructed according to Equation (13) using the calibration set. Local conformal prediction sets $\tilde{C}_{\text{loc}}$ are obtained using the local conformalisation procedure, where the support $\mathcal{X}$ is divided into 5 equal-sized bins.
@@
We then compare the average coverage and volume of the conformal prediction set $\tilde{C}$ and the local conformal prediction set $\tilde{C}_{\text{loc}}$ over 100 repetitions with two alternative conformal prediction intervals: $\tilde{C}_{q2}$, based on a quantile regression assuming a quadratic model, and $\tilde{C}_{q3}$, based on a quantile regression assuming a cubic model. See Romano et al., (2019) and Appendix E for more details about construction of prediction interval based on quantile methods. The coverage is computed by randomly generating 5,000 samples of $(Y^{L},Y^{U})$ and calculating the proportion of intervals covered by the prediction sets. The volume is calculated as the integrated Lebesgue measure of the prediction sets over the support $\mathcal{X}$.
@@
Figure 3 shows the comparison of the prediction sets from the numerical experiments. The prediction sets $\tilde{C}$ and $\tilde{C}_{\text{loc}}$ capture the bifurcation region of the conditional distribution in Model A and Model B and have smaller volume compared to the quantile regression-based prediction sets. Even when the conditional distribution is uni-modal, and the oracle prediction set is a single interval for each $x\in\mathcal{X}$, our proposed estimators still have smaller volume as demonstrated in Model C. In terms of coverage, $\tilde{C},\tilde{C}_{\text{loc}}$ and $\tilde{C}_{q3}$ appear to have coverage close to the nominal level $0.91-\alpha=$ 0.9. Notably, $\tilde{C}_{\text{loc}}$ appears to have better coverage than $\tilde{C}$ near the boundaries of the support $\mathcal{X}$.
@@ 6.1 US CPS data
We applied our method to the 2023 US Current Population Survey (CPS) dataset and construct prediction interval for their income given years of education and experience. Following Heckman et al., (2008), we construct the following variables from the dataset for individuals who report being American citizen and employed:
@@
  • Education: Following Heckman et al., (2008), we construct the years of education for each individual by converting the highest qualification they have obtained into years of education.
@@
The CPS dataset consists of 146,133 observations. There are 53,171 individuals who satisfy the demographic restrictions above. Among the individuals that satisfy the demographic criteria, 39,333 reported their income either exactly or as a band and the response rate is thus around 74%. The issue of nonresponse in surveys has been studied in the literature; see for example, Lillard et al., (1986) and Bollinger et al., (2019), where they document an increasing proportion of nonresponse of earnings in survey data and the potential bias it introduces. We will not address the nonresponse issue in this paper, but it is an important consideration in practice. Our proposed predictions could assist in the imputation of missing values in a robust manner; see Lei and Candès, (2021). In fact, allowing respondents to report their income in bands could help mitigate the non-response problem. For the 39,333 individuals who reported their incomes and satisfy the demographic restrictions, 8,819 of them reported either their gross or net income as a band, which thus amounts to roughly 22.4%. of the respondents. Figure 2 shows the proportion of individuals who reported their income in a range by years of education.
@@
In our empirical study, we first randomly reserve 20% of the data as a hold-out dataset. The remaining 80% is then further divided into training and calibration sets, with 75% of the data used for training to estimate the prediction interval $\hat{C}_{I}$ and 25% used for constructing the conformalized prediction interval $\tilde{C}_{I}$. Table 1 presents the average estimated prediction intervals across the 100 iterations. The method column indicates the way we estimate the prediction interval. “P” stands for prediction interval $\hat{C}_{I}$ in (7) and “CP” stands for conformal prediction interval $\tilde{C}_{I}$. If “M” is appended, it indicates the prediction intervals are estimated with imputed income for individuals who reported their income in an interval. It can be seen from the table that we would in general predict a higher income for individuals with more years of education.
@@
The impact of including interval-valued income data on the prediction interval is more pronounced at smaller $\alpha$ values, as evidenced by the results for $\alpha=0.1$ in the table. This is because, with a 90% coverage target, the tails of the income distribution become more critical. Given that 20% of the observations are reported in ranges and many fall into the highest category, incorporating interval-valued data shifts the right tail of the income distribution. Consequently, the prediction intervals constructed with interval-valued data are significantly different from the ones constructed with imputed income data.
Table 2 shows the coverage of the interval valued incomes in the hold-out dataset for the conformal prediction interval $\tilde{C}$ that are constructed using both interval-valued and exactly reported income data (shown in the “Censored” columns), as well as for the conformal prediction intervals constructed when the interval-valued incomes are imputed (shown in the “Imputed” columns). The significant difference in prediction intervals constructed with interval-valued data and with imputed data may be due to the fact that there are only five categories for individuals who report their income in a range. This seems somewhat coarse in terms of informing about the true income distribution. Notice that the conformal coverage using the censored model has coverage close to the nominal rate while the Imputed Coverage is almost always lower. Note that in Tables 1 an 2 when we display different level of coverage $\alpha$, this provides an empirical estimate of the $(1-\alpha)$prediction sets which can be informative about the shape of the data distribution. Notice that standard reporting of confidence intervals on parameters is meant to summarize sampling uncertainty, while here not only do these conformalized sets take into account sampling uncertainty, but they also estimate prediction sets of the underlying distributions of interest.
7 Conclusion
Model A Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Model B Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Model C Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Comparison of prediction sets in simulation study. Row 1-4 shows the observed samples $(X_{i},Y^{L}_{i},Y_{i}^{U})$ and the conformal prediction set $\tilde{C}$, the local conformal prediction set $\tilde{C}_{\text{loc}}$, the conformal prediction set constructed with quadratic quantile regression $\tilde{C}_{q2}$, and the conformal prediction set constructed with cubic quantile regression $\tilde{C}_{q3}$. Row 5 shows the integrated volume of the prediction sets, and row 6 shows the coverage of the prediction sets.
@@ 0.9 CP 10 (37,280 - 47,224) (51,494 - 64,217) (50,508 - 59,915) (64,290 - 77,790)
Supplementary Material for “Prediction Sets and Conformal Inference with Censored Outcomes”
@@ A.1 Proofs for Section 2
which shrinks nicely to $(a,b)$ (Folland,, 1999), and as a reult for some $k^{*}$, $\delta=\int_{A_{k^{*}}}f_{Y^{L},Y^{U}}>0$. Let $P_{\pi/2}$ be the one defined in Lemma A.1 and $P_{\pi/2}^{\prime}$ be the positive measure constructed with the same method for the density function $f_{Y^{L},Y^{U}}\mathbb{I}\left\{A_{k^{*}}^{c}\right\}$. For any $y\in[a,b]$, and any $\lambda>0$, let $0<\varepsilon<\delta/\lambda$, and define the following probability distribution, for any interval $B\subset\mathbb{R}$,
@@
Since $P(Y\in A)\geq P([Y^{L},Y^{U}]\subset A)$, it is clear that the outer region of the level set with $1-\alpha$ coverage of $(Y^{L},Y^{U})$ also has $1-\alpha$ coverage for $Y$. But the outer region is in general not the smallest valid prediction set.
The following lemma shows that even though our approach is conservative in terms of coverage, its conservativeness is only due to partial identification, which is necessary without assuming the mechanism of censoring.
@@
$P_{2}(A)=\int_{A_{\pi}}g_{x}\mathbb{I}\left\{\cup A^{\prime}_{m}\right\},$
@@ A.2 Proofs for Section 3
In this subsection, we first prove the uniform consistency of the proposed prediction set when the oracle prediction set is an interval under random censoring. Then the result is extended to the case allowing for fixed censoring and when oracle prediction set is a union of multiple intervals.
@@
    In the first case, we have $\hat{\tau}_{1}-\hat{\tau}_{0}\leq\tau_{1}(x)-\tau_{0}(x)$, if $E_{0,x}$ also holds, then $P(\hat{\tau}_{0},\hat{\tau}_{1};x)\leq 1-\alpha-\delta$ by Assumption 3.2, while $P_{n}(\hat{\tau}_{0},\hat{\tau}_{1};x)\geq 1-\alpha$, which implies that
    $\left\lvert P_{n}(\hat{\tau}_{0},\hat{\tau}_{1};x)-P\left(\hat{\tau}_{0},\hat{\tau}_{1};x\right)\right\rvert>\delta>\kappa.$
    Hence $E_{1,x}E_{0,x}\subset E^{c}$.
@@
For any $\varepsilon>0$, let $\delta$ be chosen as in Assumption 3.5. Let $E$ denote the event when $\sup_{C,x}\left\lvert P_{n}(C;x)-P(C;x)\right\rvert\leq\psi_{n}$. For sufficiently large $n$, $\psi_{n}\leq\frac{\delta}{2}$. For any $x\in\mathcal{X}$, if $\mu(\hat{C}(x)\bigtriangleup C^{*}(x))>\varepsilon$ and $\mu(\hat{C}(x))\leq\mu(C^{*}(x))$, then $P(\hat{C}(x);x)\leq 1-\alpha-\delta$, and $P_{n}(\hat{C}(x);x)\geq 1-\alpha-\psi_{n}$. Hence $\left\lvert P_{n}(\hat{C};x)-P(\hat{C};x)\right\rvert>\frac{\delta}{2}\geq\psi_{n}$. On the other hand, if $\mu(\hat{C}(x)\bigtriangleup C^{*}(x))>\varepsilon$ and $\mu(\hat{C}(x))>\mu(C^{*}(x))$, then $P_{n}(C^{*}(x);x)<1-\alpha-\psi_{n}$ while $P(C^{*}(x);x)\geq 1-\alpha$ and $\left\lvert P_{n}(C^{*}(x);x)-P(C^{*}(x);x)\right\rvert>\psi_{n}$. That is $P(\sup_{x}\mu(\hat{C}(x)\bigtriangleup C^{*}(x))>\varepsilon)\leq P(E^{c})\to 0$. ∎
@@ A.3 Proofs for results in Section 4
Let $s_{i}=s(Y^{L}_{i},Y^{U}_{i},X_{i};\hat{C})$, then $[Y^{L}_{n+1},Y^{U}_{n+1}]\subset\tilde{C}(X_{n+1})$ if and only if $s_{n+1}\leq\vartheta_{1-\alpha}$. Since $s_{i}$’s are exchangeable for $i\in\mathcal{I}_{2}$, $P\left(s_{n+1}\leq\vartheta_{1-\alpha}\right)\geq 1-\alpha$ by Lemma 2 in Romano et al., (2019). ∎
@@
Fix an $\varepsilon>0$, on the event $\sup_{x}\mu\left(\hat{C}(x)\bigtriangleup C^{*}(x)\right)<\frac{\varepsilon}{2}$, we know that
$\left\lvert s(y^{l},y^{u},x;\hat{C})-s(y^{l},y^{u},x;C^{*})\right\rvert<\frac{\varepsilon}{2},$
for all $y^{l},y^{u}$ and $x$. Let $s_{[j]},s^{*}_{[j]}$ be the corresponding order statistics for $s_{j}$ and $s^{*}_{j},j\in\mathcal{I}_{2}$. Let $s_{[l]}=s_{[l]}$ when $l$ is an integer and $s_{[l]}=s_{[\lfloor l\rfloor+1]}$ when $l$ is not an integer. We have $s_{[(n_{2}+1)(1-\alpha)]}>s^{*}_{[(n_{2}+1)(1-\alpha)]}-\varepsilon$ provided that $s^{*}_{[(n_{2}+1)(1-\alpha)]}-s^{*}_{[(n_{2}+1)(1-\alpha)-1]}<\frac{\varepsilon}{2}$ which happens with high probability for sufficiently large $n_{2}$. Similarly we have $s_{[(n_{2}+1)(1-\alpha)]}<s^{*}_{[(n_{2}+1)(1-\alpha)]}+\varepsilon$. Thanks to the consistency of sample quantile, we have that $s^{*}_{[(n_{2}+1)(1-\alpha)]}\to_{p}Q_{1-\alpha}(s^{*})$ as $n_{2}\to\infty$. By the design of the score function, $[y^{L},y^{U}]\subset C(x)$ if and only if $s(y^{L},y^{U},x;C(x))\leq 0$. Since $Q_{1-\alpha}(s^{*})=\inf\left\{z:P(s^{*}\leq z)\geq 1-\alpha\right\}$, combined with
$P(s^{*}\leq 0)=P\left([Y^{L},Y^{U}]\subset C^{*}\right)\geq 1-\alpha,$
and for any $z<0$, $P(s^{*}\leq z)<1-\alpha$ by Assumption 3.2, we have $Q_{1-\alpha}(s^{*})=0$. Hence $\vartheta_{1-\alpha}=o_{p}(1)$. The proof for the local version is similar and omitted. ∎